OpenStudy (anonymous):

Find the minimum, maximum, range, mean, and median for the data set. 12, 13, 13, 17, 18, 19, 27 A. minimum: 12 maximum: 27 range: 15 mean: 17 median: 13 B. minimum: 27 maximum: 12 range: 15 mean: 17 median: 17 C. minimum: 12 maximum: 27 range: 15 mean: 17 median: 17 D. minimum: 12 maximum: 27 range: 17 mean: 15 median: 17

OpenStudy (anonymous):

@Nurali

OpenStudy (inowalst):

First, whats the min?

OpenStudy (anonymous):

I Think its A) Idk..

OpenStudy (inowalst):

Do you know how to find range?

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

No >_<

OpenStudy (anonymous):

Is it A) ?

OpenStudy (anonymous):

Wait No B..

OpenStudy (bohotness):

wahat is in the middle ?

OpenStudy (inowalst):

The range of a set of data is the difference between the highest and lowest values in the set.

OpenStudy (inowalst):

The mean is the average of the numbers. To find the mean, add up all the numbers, then divide by how many numbers there are.

OpenStudy (nurali):

12, 13, 13, 17, 18, 19, 27 Minimum=12 Maximum=27 Range=Maximum-Minimum=27-12=15 Mean=The mean is the usual average, so 12+13+13+17+18+19+27/7=17 Mode=The mode is the number that is repeated more often than any other, so 13 is the mode.

OpenStudy (anonymous):

C* Im Sorry My typoes Today

OpenStudy (anonymous):

..

OpenStudy (inowalst):

To find median, order the numbers from least to greatest, then the middle number is the median. If you're left with two in the middle, add them, then divide them by 2.

OpenStudy (anonymous):

Smallest Number is 12 Bigest number is 27

OpenStudy (anonymous):

Range = Maximum - Minimum = 27 - 12 = 15

OpenStudy (anonymous):

Mean = (12+13+13+17+18+19+27)/7 = 119/7 = 17

OpenStudy (anonymous):

Median = The middle value (If there are 7 data points, then you want the 4th-ranked value) = 17?

OpenStudy (inowalst):

@Nurali They ask for median, not mode.

OpenStudy (nurali):

Median=The median is the middle value, so 17